OpenStudy (anonymous):

Hello, I have an algebra word pro blem that I dont know how to set up so I can solve it. Can someone help me please?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

ask it;

OpenStudy (amistre64):

2 people working together?

OpenStudy (anonymous):

Ok here is the word problem: The sum of both digits, of either of two two-digit numbers, in whatever order the digits are written, is 9. The square of either of the digits of either number, minus the product of both digits, plus the square of the other digit is the number 21. The numbers are ?

OpenStudy (amistre64):

a+b = 9 a^2 - ab +b^2 = 21

OpenStudy (amistre64):

b = 9-a so subit in

OpenStudy (anonymous):

OMG. How do I start that?

OpenStudy (amistre64):

a^2 - a(9-a) +(9-a)^2 = 21

OpenStudy (amistre64):

a^2 -9a +a^2 +81 +a^2 -18a = 21

OpenStudy (amistre64):

3a^2 -27a +60 = 0

OpenStudy (anonymous):

How are u coming up with this?

OpenStudy (amistre64):

3(a^2 -9a +20) = 0 a = -(-5,-4) = 5 and 4

OpenStudy (amistre64):

magic fingers lol

OpenStudy (anonymous):

Well I need those fingers!

OpenStudy (amistre64):

25 + 16 - 36 = 21? 61 36 --- 21

OpenStudy (amistre64):

i took what they gave me and determined a suitable value for b, so that i could put the whole thing in one variable and solve it for that variable

OpenStudy (anonymous):

idk.